# Coupling Generative Modeling and an Autoencoder with the Causal Bridge

### v1

@@ 1 Introduction
Estimating the causal effect of a treatment on an outcome is crucial in various domains, but the presence of unobserved confounders can hinder accurate inference (pearl2016causal, ; Rubin_potential_outcomes, ). Traditional methods often rely on strong assumptions, such as the absence of unobserved confounders (rubin1974estimating, ; unconfounding, ). Other approaches have assumed available instrumental variables with specific properties (Rubin_IV, ; Gretton_IV, ; DeepIV, ). However, these assumptions may not always hold in practice, leading to biased estimates of causal effects.
A promising approach to address this challenge is the use of proxy variables (frost1979proxy, ; miao2018identifying, ; VAE_causal, ), also known as negative control variables. These are variables that are affected by the unobserved confounder, but do not necessarily directly influence the treatment or outcome. Consequently, leveraging information from these proxies, we can gain insight into the underlying causal mechanisms and potentially mitigate the bias caused by unobserved confounders.
Recent work has introduced the concept of a causal bridge function, which uses two sets of proxy variables, one related to the treatment and the other to the outcome, to estimate causal effects (miao2018identifying, ; Miao2020, ; gretton_causal, ). This approach has shown promising results, but there is still room for improvement in terms of both theoretical understanding and practical implementation.
This paper builds on the causal bridge framework and makes several key advances. We provide a refined theoretical analysis, clarifying the assumptions and conditions under which the causal bridge function yields accurate causal effect estimates. Furthermore, we introduce a novel learning approach that leverages the power of generative models to enhance the estimation of the causal bridge. Our approach enables the sharing of statistical strength between observed variables, leading to more robust and accurate causal inference. Finally, we extend the causal bridge framework to handle survival outcomes, a common type of data in biomedical applications. In summary:
@@ 2 Related Work
Traditional approaches to dealing with unobserved confounders often involve sensitivity analysis (Sensitivity, ; sensitivity_no_assumptions, ), which assesses the robustness of causal effect estimates to different assumptions about the unobserved confounder (binary_latent, ; Sensitivity_Robins, ; imbens_confounding, ; Sensitivity, ; kallus_confounding, ). Another common approach is the use of instrumental variables (IVs), variables that influence the treatment but are independent of the unobserved confounder (Imben_IV, ; Rubin_IV, ; DeepIV, ; Gretton_IV, ). However, finding valid IVs can be challenging in practice.
More recently, there has been growing interest in using proxy variables to address unobserved confounding. Early work focused on categorical data and unobserved confounders (kuroki2014measurement, ), showing that causal effects can be estimated under certain conditions. The concept of causal bridge function (miao2018identifying, ; Miao2020, ; gretton_causal, ) extends this idea to more general settings, using two sets of proxies to estimate causal effects. These studies have provided valuable theoretical insights and practical tools for causal inference with proxy variables.
Deep learning has also been increasingly applied to causal inference, enabling the development of more flexible models. For example, deep-generative models have been used to improve IV analysis (DeepIV, ) and to address unobserved confounding in various settings (VAE_causal, ; GANITE, ; assaad, ). Our work builds on these advances, leveraging generative models to enhance the estimation of the causal bridge function.
@@ 3 Background on the Causal Bridge
Let $X$ and $Y$ represent a treatment (intervention) and outcome, respectively. In the examples we consider here, $Y$ will be continuous and $X$ could be real or categorical (with binary being an important special case). It is assumed that $X$ and $Y$ are both dependent on an unobserved confounder $U$, as illustrated in Figure 1. As discussed in (kuroki2014measurement, ; Pearl_book, ; miao2018identifying, ; Miao2020, ; gretton_causal, ), to perform inference in the presence of the unobserved $U$, control (proxy) measurements are assumed to be available, which are also dependent on $U$. Specifically, $Z$ is a treatment control variable, and $W$ is an outcome control variable. As shown in Figure 1, the outcome may depend on $W$ and the treatment may depend on $Z$. We assume for simplicity that there are no additional covariates, but such could be included if available, as discussed in (Miao2020, ; gretton_causal, ). We denote the domains of $(Y,X,Z,W)$ as $(\mathcal{Y},\mathcal{X},\mathcal{Z},\mathcal{W})$, respectively.
For the graphical model in Figure 1, we make the following assumptions, which are consistent with those in (miao2018identifying, ; Miao2020, ; gretton_causal, ):
@@
$Z\perp\!\!\!\!\perp Y|(U,X)\hskip 5.69054pt{\rm and}\hskip 5.69054ptZ\perp\!\!\!\!\perp W|(U,X)$
@@
We now introduce a bridge function $b(W,x)$ (Miao2020, ; gretton_causal, ; cui2024semiparametric, ).
Theorem 1 cui2024semiparametric  If there is a solution to the Fredholm integral equation
@@
The form of Theorem 1 was presented and proven in cui2024semiparametric , and requires multiple additional assumptions to ensure that there is a solution to (1) (miao2018identifying, ; gretton_causal, ), and importantly, completeness, described below.
@@
Theorem 2 cui2024semiparametric Under A1-A4, and other technical conditions, there exists a solution to (1).
The proof and further details concerning Theorem 2 are presented in cui2024semiparametric . The above assumptions (and other technical requirements) are relatively abstract; therefore, in Section 4 below we explore the properties that must hold for the existence of a solution to (1). The insights from this analysis will motivate our modeling approach in Section 6.
@@ 4 Implied Properties of the Bridge Function
Note that (7) is by definition and is not an assumption. Moreover, condition (6) requires (1) to hold. This interpretation of the bridge function $b(W,x)$ as matching $b_{0}(W,x,z)$ in expectation rather than a pointwise (distributional) match is consistent with relaxed identification frameworks discussed in the literature of nonparametric instrumental variables (newey2003instrumental, ; carrasco2007linear, ), where moment conditions or expectations replace the exact operator equalities.
@@
Examining the statement of Theorem 3, one may be tempted to suggest proxies $Z$ that are independent of $U$, but the completeness assumption (A4) concerning $\mathbb{E}[g(U)|Z,x]$ disallows it. Instead, Theorem 3 states that $W$ should be a low-noise representation of $U$, in the sense discussed in kuroki2014measurement concerning proxies $(W,Z)$ as “noisy” measurements of $U$.
@@
The complete statement of Corollary 1 and its proof are provided in Appendix B.2. While the assumption of an invertible $\Psi(U)$ may seem strong, this same assumption was made previously for a similar setup in kuroki2014measurement , where discrete observations and proxies were considered (see Eq. (4) in kuroki2014measurement ).
@@
Figure 2 shows the results for the relative approximation error $r(\eta)=\mathbb{E}_{Z\sim p(Z|x)}[\eta/|\mathbb{E}[Y|x,Z]|]$ vs. $I(U;Z|W,x)$ both averaged over $X$, for $\sigma_{U}=10$, $\sigma_{Z},\sigma_{W}=\{0.1,0.25,0.5,0.75,1\}$, and $\sigma_{X}=0.1$, from which we see that $\eta$ increases with $I(U;Z|W,x)$ and $\sigma_{W}$, consistent with Theorem 3. We provide more details and results varying $\sigma_{X}$ in Appendix D. Interestingly, when $\sigma_{X}=\sigma_{Z}$ and the SEM coefficients are 1, both $\eta=0$ and $I(U;Z|W,x)=0$, which is a special case formalized in Lemma 1 in Appendix C.
@@ 5 Generative Model for the Bridge Function
We seek to model the form of the bridge as in (10). From that perspective, a generative model can be constructed to draw the samples of $U$ needed for implementing (1). In this context, and using (10)), we consider $u_{j}\sim p(U|w_{j},x)$ with $w_{j}\sim p(W|x,z)$. One may consider learning a function $h(W,x,\epsilon)$, such that $u_{j}=h(w_{j},x,\epsilon_{j})$, where $w_{j}\sim p(W|x,z)$ and (for example) $\epsilon_{j}\sim\mathcal{N}(0,I)$, where $I$ is the identity matrix and hence $\epsilon$ is drawn from isotropic Gaussian noise of the chosen dimension. Such a generative model $h(W,x,\epsilon)$ has been widely considered, e.g., in generative adversarial networks (GANs) and its generalizations (GAN, ; fGAN, ; wgan, ; infogan, ; cGAN, ). In practice, using Assumption 5, we may model
@@ 6 Learning Setup
Assume that we have access to a set of data $\mathcal{D}_{1}=\{(x_{i},z_{i},w_{i})\}_{i=1,M}$ from which a generative model can be learned to draw samples from $p(W|x,z)$. The details of this model depend on the data characteristics, and more details are presented when discussing the experiments in Section 7. Note that we need not model $p(W|x,z)$, rather in general, we seek to model the capacity to generate samples from this conditional distribution, e.g., using a conditional GAN (cGAN, ).
@@
The objective implied by $\mathcal{L}_{\theta_{Y}}$ involves two steps: ($i$) develop a generative model to draw samples from $p(W|x,z)$ using ${\cal D}_{1}$, and ($ii$) use this model within the minimization of $\mathcal{L}_{\theta_{Y}}$ for $\theta_{Y}$ to approximate the expectation wrt $W$. However, note that these two steps are followed in sequence, which should be distinguished from the iterative and alternating two-step approach developed in gretton_causal .
@@
Autoencoder for the Treatment and its Control Analogous to the aforementioned model for $Y$, we can model $\mathbb{E}(X|U,z)=g_{\theta_{X}}(U,z)$ and $\mathbb{E}(Z|U)=g_{\theta_{Z}}(U)$. In the context of an autoencoder, we assume that $u_{j}=h_{\theta_{U}}(w_{j},x,\epsilon_{j})$ with $w_{j}\sim p(W|x_{i},z_{i})$ constitutes a means of encoding observations $(x_{i},z_{i})$ into a latent feature space represented by conditional samples $u_{j}|(x_{i},z_{i})$. Connected to the approximation considered in Proposition 2, we assume that the expectations of $g_{\theta_{X}}(U,z)$ and $g_{\theta_{Z}}(U)$ wrt $p(U|W,x,z)$ may be replaced by expectations wrt $p(U|W,x)$ with $W\sim p(W|x,z)$. This implies the same conditions on $p(U|W,x,z)$, e.g., that the expectation of $p(U|W,x,z)$ depends only on $(W,x)$, and that $g_{\theta_{X}}(U,z)$ and $g_{\theta_{Z}}(U)$ have the same class of dependence on $U$ as $g_{\theta_{Y}}(x,W,U)$, e.g., they could be linear in $U$. However, although these functions may be linear in $U$, $U$ itself may be nonlinearly related to $W$ (see Corollary 1). From this angle, we consider the additional losses:
$\begin{aligned} \mathcal{L}_{\theta_{X}}&=\textstyle{\sum}_{i=1}^{N}\big(x_{i}-\mathbb{E}_{p(W|x_{i},z_{i})}\mathbb{E}_{p(\epsilon)}[g_{\theta_{X}}(h_{\theta_{U}}(W,x,\epsilon),z_{i})]\big)^{2}\\ \mathcal{L}_{\theta_{Z}}&=\textstyle{\sum}_{i=1}^{N}\big(z_{i}-\mathbb{E}_{p(W|x_{i},z_{i})}\mathbb{E}_{p(\epsilon)}[g_{\theta_{Z}}(h_{\theta_{U}}(W,x,\epsilon))]\big)^{2},\end{aligned}$ (13)
@@
Connection to the Causal Effects VAE (CEVAE)  The composite model, employing the cumulative loss function $\mathcal{L}_{\theta_{Y}}+\mathcal{L}_{\theta_{X}}+\mathcal{L}_{\theta_{Z}}$, shares characteristics and motivation with the CEVAE developed in VAE_causal . We note that recent work by CEVAE_critical ; gretton_causal has highlighted limitations of the CEVAE. Some of the challenges with existing CEVAE research are related to the difficulty of modeling posteriors like $p(U|x,z)$.
@@
The extension of our approach to a CEVAE-type setup is relatively straightforward, using our definition of $p(U|x,z)$ within the CEVAE. Importantly, in such a setting, the CEVAE models $(X,Z)$, and $Y$ is handled via the Fredholm equation for the bridge. The main difference between a CEVAE version of our approach is the inclusion of a prior $p(U)$ and the inclusion of a KL term between $p(U)$ and $p(U|x,z)$. Nevertheless, we did implement such a modified CEVAE formulation, in addition to the simpler autoencoder setup discussed above. We found that the KL term added significant difficulty and undermined the reliability of CEVAE-based predictions. We do not show the CEVAE results in Section 7, because they were numerically unstable and sensitive to the way $p(U)$ was set (like shown in gretton_causal ). In contrast, we found that our approach trained well and yielded reliable results.
Bridge for Survival (Time to Event) Outcomes So far we have considered continuous outcomes $Y$ with standard squared error loss $\mathcal{L}_{\theta_{Y}}$. We now consider survival outcomes, which are of special interest in a wide range of scenarios where causal inference is used in practice, constituting a novel application of the causal-bridge framework. Specifically, we consider outcomes of the form $(Y,E)$, where $Y$ is the observed time $Y=\min(T,C)$, $T$ is the time at which the event of interest occurs, $C$ is the follow-up time, and $E$ is the observed-event indicator. If for a given sample, $y=t<c$ it is said that the event of interest is observed and $e=1$, otherwise, $y=c<t$, $e=0$ and the event is right-censored. Here we assume that censoring is not informative, i.e., $T\perp\!\!\!\!\perp C|X$. Extensions to other forms of censoring are possible within our framework, but left as future work. Unlike for continuous outcomes, we are not interested in modeling the (expected) value of $Y$ through $\mathbb{E}[Y|do(X=x)]$. Instead, we are interested in $\mathbb{E}[\lambda|do(X=x)]$, i.e., the risk function defined as the contribution of observed covariates on a baseline hazard function, i.e., $\lambda(Y|W,x)=\lambda_{0}(t)\exp(b(W,x))$, where $\lambda_{0}(t)$ is the baseline hazard and $\exp(b(W,x))$ is the risk function, conveniently written in terms of a bridge function. The casual estimate of interest for binary treatments, $X=\{0,1\}$ is the hazards ratio (HR) defined as $\lambda(Y|W,X=1)/\lambda(Y|W,X=0)=\exp(b(W,X=1))/\exp(b(W,X=0))$. Optimizing (1) wrt the hazards function is achieved by maximizing the partial likelihood of the model similar to the Cox proportional hazard model cox1972regression using
@@
Recently, (ying2022proximal, ) proposed an approach that also modeled the hazard function using the bridge function. However, they do not make the proportional hazard assumption like we do. Rather, they impose a rigid form for the bridge function (Eq. (31) in (ying2022proximal, )). That work considered experiments on real data (SUPPORT), as we do. However, in (ying2022proximal, ) the ground truth for the estimate of the causal effect is not available, while in our experiments in Section 7 we compare to an RCT.

### v2

@@ 1 Introduction
Estimating the causal effect of a treatment on an outcome is crucial in various domains, but the presence of unobserved confounders can hinder accurate inference (Pearl et al., 2016; Rubin, 2005). Traditional methods often rely on strong assumptions, such as the absence of unobserved confounders (Rubin, 1974; Rosenbaum, 1984). Other approaches have assumed available instrumental variables with specific properties (Angrist et al., 1996; Xu et al., 2021a; Hartford et al., 2017). However, these assumptions may not always hold in practice, leading to biased estimates of causal effects.
A promising approach to address this challenge is the use of proxy variables (Frost, 1979; Miao et al., 2018; Louizos et al., 2017), also known as negative control variables. These are variables that are affected by the unobserved confounder, but do not necessarily directly influence the treatment or outcome. Consequently, leveraging information from these proxies, we can gain insight into the underlying causal mechanisms and potentially mitigate the bias caused by unobserved confounders.
Recent work has introduced the concept of a causal bridge function, which uses two sets of proxy variables, one related to the treatment and the other to the outcome, to estimate causal effects (Miao et al., 2018, 2024; Xu et al., 2021b). This approach has shown promising results, but there is still room for improvement in terms of both theoretical understanding and practical implementation.
This paper builds on the causal bridge framework and makes several key advances. We provide a refined theoretical analysis, clarifying the assumptions and conditions under which the causal bridge function yields accurate causal effect estimates. Furthermore, we introduce a novel learning approach that leverages the power of generative models to enhance the estimation of the causal bridge. Our approach enables the sharing of statistical strength between observed variables, leading to more robust and accurate causal inference. Finally, we extend the causal bridge framework to handle survival outcomes, a common type of data in biomedical applications.
In summary:
@@ 2 Related Work
Traditional approaches to dealing with unobserved confounders often involve sensitivity analysis (Franks et al., 2020; Ding and VanderWeele, 2016), which assesses the robustness of causal effect estimates to different assumptions about the unobserved confounder (Rosenbaum and Rubin, 1983a; Robins et al., 2000; Imbens, 2003; Franks et al., 2020; Kallus et al., 2019). Another common approach is the use of instrumental variables (IVs), variables that influence the treatment but are independent of the unobserved confounder (Angrist and Imbens, 1995; Angrist et al., 1996; Hartford et al., 2017; Xu et al., 2021a). However, finding valid IVs can be challenging in practice.
More recently, there has been growing interest in using proxy variables to address unobserved confounding. Early work focused on categorical data and unobserved confounders (Kuroki and Pearl, 2014), showing that causal effects can be estimated under certain conditions. The concept of causal bridge function (Miao et al., 2018, 2024; Xu et al., 2021b) extends this idea to more general settings, using two sets of proxies to estimate causal effects. These studies have provided valuable theoretical insights and practical tools for causal inference with proxy variables.
Deep learning has also been increasingly applied to causal inference, enabling the development of more flexible models. For example, deep-generative models have been used to improve IV analysis (Hartford et al., 2017) and to address unobserved confounding in various settings (Louizos et al., 2017; Yoon et al., 2018; Assaad et al., 2021). Our work builds on these advances, leveraging generative models to enhance the estimation of the causal bridge function.
@@ 3 Background on the Causal Bridge
Let $X$ and $Y$ represent a treatment (intervention) and outcome, respectively. In the examples we consider here, $Y$ will be continuous and $X$ could be real or categorical (with binary being an important special case). It is assumed that $X$ and $Y$ are both dependent on an unobserved confounder $U$, as illustrated in Figure 1. As discussed in (Kuroki and Pearl, 2014; Pearl, 2009; Miao et al., 2018, 2024; Xu et al., 2021b), to perform inference in the presence of the unobserved $U$, control (proxy) measurements are assumed to be available, which are also dependent on $U$. Specifically, $Z$ is a treatment control variable, and $W$ is an outcome control variable. As shown in Figure 1, the outcome may depend on $W$ and the treatment may depend on $Z$. We assume for simplicity that there are no additional covariates, but such could be included if available, as discussed in (Miao et al., 2024; Xu et al., 2021b). We denote the domains of $(Y,X,Z,W)$ as $(\mathcal{Y},\mathcal{X},\mathcal{Z},\mathcal{W})$, respectively.
For the graphical model in Figure 1, we make the following assumptions, which are consistent with those in (Miao et al., 2018, 2024; Xu et al., 2021b):
@@
$Z\perp\!\!\!\!\perp Y|(U,X)\hskip 5.69054pt{\rm and}\hskip 5.69054ptZ\perp\!\!\!\!\perp W|U$
@@
We now introduce a bridge function $b(W,x)$ (Miao et al., 2024; Xu et al., 2021b; Cui et al., 2024).
Theorem 1 Cui et al. (2024)  If there is a solution to the Fredholm integral equation
@@
The form of Theorem 1 was presented and proven in Cui et al. (2024), and requires multiple additional assumptions to ensure that there is a solution to (1) (Miao et al., 2018; Xu et al., 2021b), and importantly, completeness, described below.
@@
Theorem 2 Cui et al. (2024) Under A1-A4, and other technical conditions, there exists a solution to (1).
The proof and further details concerning Theorem 2 are presented in Cui et al. (2024). The above assumptions (and other technical requirements) are relatively abstract; therefore, in Section 4 below we explore the properties that must hold for the existence of a solution to (1). The insights from this analysis will motivate our modeling approach in Section 6.
@@ 4 Implied Properties of the Bridge Function
Note that (7) is by definition and is not an assumption. Moreover, condition (6) is required for (1) to hold. This interpretation of the bridge function $b(W,x)$ as matching $b_{0}(W,x,z)$ in expectation rather than a pointwise (distributional) match is consistent with relaxed identification frameworks discussed in the literature of nonparametric instrumental variables (Newey and Powell, 2003; Carrasco et al., 2007), where moment conditions or expectations replace the exact operator equalities.
@@
Examining the statement of Theorem 3, one may be tempted to suggest proxies $Z$ that are independent of $U$, but the completeness assumption (A4) concerning $\mathbb{E}[g(U)|Z,x]$ disallows it. Instead, Theorem 3 states that $W$ should be a low-noise representation of $U$, in the sense discussed in Kuroki and Pearl (2014) concerning proxies $(W,Z)$ as “noisy” measurements of $U$.
@@
The complete statement of Corollary 1 and its proof are provided in Appendix B.2. While the assumption of an invertible $\Psi(U)$ may seem strong, this same assumption was made previously for a similar setup in Kuroki and Pearl (2014), where discrete observations and proxies were considered (see Eq. (4) in Kuroki and Pearl (2014)).
@@
Figure 2 shows the results for the relative approximation error $r(\eta)=\mathbb{E}_{Z\sim p(Z|x)}[\eta/|\mathbb{E}[Y|x,Z]|]$ vs. $I(U;Z|W,x)$ both averaged over $X$, for $\sigma_{U}=10$, $\sigma_{Z},\sigma_{W}=\{0.1,0.25,0.5,0.75,1\}$, and $\sigma_{X}=0.1$, from which we see that $\eta$ increases with $I(U;Z|W,x)$ and $\sigma_{W}$, consistent with Theorem 3; $\eta$ is defined in (8). We provide more details and results varying $\sigma_{X}$ in Appendix D. Interestingly, when $\sigma_{X}=\sigma_{Z}$ and the SEM coefficients are 1, both $\eta=0$ and $I(U;Z|W,x)=0$, which is a special case formalized in Lemma 1 in Appendix C.
@@ 5 Generative Model for the Bridge Function
We seek to model the form of the bridge as in (10). From that perspective, a generative model can be constructed to draw the samples of $U$ needed for implementing (1). In this context, and using (10)), we consider $u_{j}\sim p(U|w_{j},x)$ with $w_{j}\sim p(W|x,z)$. One may consider learning a function $h(W,x,\epsilon)$, such that $u_{j}=h(w_{j},x,\epsilon_{j})$, where $w_{j}\sim p(W|x,z)$ and (for example) $\epsilon_{j}\sim\mathcal{N}(0,I)$, where $I$ is the identity matrix and hence $\epsilon$ is drawn from isotropic Gaussian noise of the chosen dimension. Such a generative model $h(W,x,\epsilon)$ has been widely considered, e.g., in generative adversarial networks (GANs) and its generalizations (Goodfellow et al., 2014; Nowozin et al., 2016; Arjovsky et al., 2017; Chen et al., 2016; Mirza and Osindero, 2014). In practice, using Assumption 5, we may model
@@ 6 Learning Setup
Assume that we have access to a set of data $\mathcal{D}_{1}=\{(x_{i},z_{i},w_{i})\}_{i=1,M}$ from which a generative model can be learned to draw samples from $p(W|x,z)$. The details of this model depend on the data characteristics, and more details are presented when discussing the experiments in Section 7. Note that we need not model $p(W|x,z)$, rather in general, we seek to model the capacity to generate samples from this conditional distribution, e.g., using a conditional GAN (Mirza and Osindero, 2014).
@@
The objective implied by $\mathcal{L}_{\theta_{Y}}$ involves two steps: ($i$) develop a generative model to draw samples from $p(W|x,z)$ using ${\cal D}_{1}$, and ($ii$) use this model within the minimization of $\mathcal{L}_{\theta_{Y}}$ for $\theta_{Y}$ to approximate the expectation wrt $W$. However, note that these two steps are followed in sequence, which should be distinguished from the iterative and alternating two-step approach developed in Xu et al. (2021b).
@@
Autoencoder for the Treatment and its Control Analogous to the aforementioned model for $Y$, we can model $\mathbb{E}(X|U,z)=g_{\theta_{X}}(U,z)$ and $\mathbb{E}(Z|U)=g_{\theta_{Z}}(U)$. In the context of an autoencoder, we assume that $u_{j}=h_{\theta_{U}}(w_{j},x_{i},\epsilon_{j})$ with $w_{j}\sim p(W|x_{i},z_{i})$ constitutes a means of encoding observations $(x_{i},z_{i})$ into a latent feature space represented by conditional samples $u_{j}|(x_{i},z_{i})$. Connected to the approximation considered in Proposition 2, we assume that the expectations of $g_{\theta_{X}}(U,z)$ and $g_{\theta_{Z}}(U)$ wrt $p(U|W,x,z)$ may be replaced by expectations wrt $p(U|W,x)$ with $W\sim p(W|x,z)$. This implies the same conditions on $p(U|W,x,z)$, e.g., that the expectation of $p(U|W,x,z)$ depends only on $(W,x)$, and that $g_{\theta_{X}}(U,z)$ and $g_{\theta_{Z}}(U)$ have the same class of dependence on $U$ as $g_{\theta_{Y}}(x,W,U)$, e.g., they could be linear in $U$. However, although these functions may be linear in $U$, $U$ itself may be nonlinearly related to $W$ (see Corollary 1). From this angle, we consider the additional losses:
$\begin{aligned} \mathcal{L}_{\theta_{X}}&=\textstyle{\sum}_{i=1}^{N}\big(x_{i}-\mathbb{E}_{p(W|x_{i},z_{i})}\mathbb{E}_{p(\epsilon)}[g_{\theta_{X}}(h_{\theta_{U}}(W,x_{i},\epsilon),z_{i})]\big)^{2}\\ \mathcal{L}_{\theta_{Z}}&=\textstyle{\sum}_{i=1}^{N}\big(z_{i}-\mathbb{E}_{p(W|x_{i},z_{i})}\mathbb{E}_{p(\epsilon)}[g_{\theta_{Z}}(h_{\theta_{U}}(W,x_{i},\epsilon))]\big)^{2},\end{aligned}$ (13)
@@
Connection to the Causal Effects VAE (CEVAE)  The composite model, employing the cumulative loss function $\mathcal{L}_{\theta_{Y}}+\mathcal{L}_{\theta_{X}}+\mathcal{L}_{\theta_{Z}}$, shares characteristics and motivation with the CEVAE developed in Louizos et al. (2017). We note that recent work by Rissanen and Marttinen (2021); Xu et al. (2021b) has highlighted limitations of the CEVAE. Some of the challenges with existing CEVAE research are related to the difficulty of modeling posteriors like $p(U|x,z)$.
@@
The extension of our approach to a CEVAE-type setup is relatively straightforward, using our definition of $p(U|x,z)$ within the CEVAE. Importantly, in such a setting, the CEVAE models $(X,Z)$, and $Y$ is handled via the Fredholm equation for the bridge. The main difference between a CEVAE version of our approach is the inclusion of a prior $p(U)$ and the inclusion of a KL term between $p(U)$ and $p(U|x,z)$. Nevertheless, we did implement such a modified CEVAE formulation, in addition to the simpler autoencoder setup discussed above. We found that the KL term added significant difficulty and undermined the reliability of CEVAE-based predictions. We do not show the CEVAE results in Section 7, because they were numerically unstable and sensitive to the way $p(U)$ was set (like shown in Xu et al. (2021b)). In contrast, we found that our approach trained well and yielded reliable results.
Bridge for Survival (Time to Event) Outcomes So far we have considered continuous outcomes $Y$ with standard squared error loss $\mathcal{L}_{\theta_{Y}}$. We now consider survival outcomes, which are of special interest in a wide range of scenarios where causal inference is used in practice, constituting a novel application of the causal-bridge framework. Specifically, we consider outcomes of the form $(Y,E)$, where $Y$ is the observed time $Y=\min(T,C)$, $T$ is the time at which the event of interest occurs, $C$ is the follow-up time, and $E$ is the observed-event indicator. If for a given sample, $y=t<c$ it is said that the event of interest is observed and $e=1$, otherwise, $y=c<t$, $e=0$ and the event is right-censored. Here we assume that censoring is not informative, i.e., $T\perp\!\!\!\!\perp C|X$. Extensions to other forms of censoring are possible within our framework, but left as future work. Unlike for continuous outcomes, we are not interested in modeling the (expected) value of $Y$ through $\mathbb{E}[Y|do(X=x)]$. Instead, we are interested in $\mathbb{E}[\lambda|do(X=x)]$, i.e., the risk function defined as the contribution of observed covariates on a baseline hazard function, i.e., $\lambda(Y|W,x)=\lambda_{0}(t)\exp(b(W,x))$, where $\lambda_{0}(t)$ is the baseline hazard and $\exp(b(W,x))$ is the risk function, conveniently written in terms of a bridge function. The casual estimate of interest for binary treatments, $X=\{0,1\}$ is the hazards ratio (HR) defined as $\lambda(Y|W,X=1)/\lambda(Y|W,X=0)=\exp(b(W,X=1))/\exp(b(W,X=0))$. Optimizing (1) wrt the hazards function is achieved by maximizing the partial likelihood of the model similar to the Cox proportional hazard model Cox (1972) using
@@
Recently, (Ying et al., 2022) proposed an approach that also modeled the hazard function using the bridge function. However, they do not make the proportional hazard assumption like we do. Rather, they impose a rigid form for the bridge function (Eq. (31) in (Ying et al., 2022)). That work considered experiments on real data (SUPPORT), as we do. However, in (Ying et al., 2022) the ground truth for the estimate of the causal effect is not available, while in our experiments in Section 7 we compare to an RCT.